You own an investment that promises to pay $100,000 thirty years from today. a. What is the present value of this security at a discount rate of 6%? 8%? 10%? b. What is meant by the phrase discounting? What is a discount rate? c. What is the relationship between present values and interest rates?
Answer #1
a] PV at 6% = 100000/1.06^30 = $     17,411.01
PV at 8% = 100000/1.08^30 = $       9,937.73
PV at 10% = 100000/1.10^30 = $       5,730.86
b] Discounting is the process of finding out the present
value of sums receivable/payable in future. In simpler
terms, it converts future sums in current dollar terms.
Discount rate is the cost of money related to time, which
provides for the following:
*Compensation for foregoing current consumption
*Income that can be earned using the money at hand
*Compensation for inflation for deferring cash flow
*Risk in deferring cash flow
c] PVs and interest rates have an inverse relationship.
That is, when interest rates go up, the PVs go down and
vice versa.